OpenStudy (anonymous):

3sec^2x-6=8tanx. Solve for x

OpenStudy (mertsj):

\[3\sec ^2x-6=8\tan \tan x\]

OpenStudy (mertsj):

Is that it?

OpenStudy (mertsj):

(With only 1 tan of course)

OpenStudy (anonymous):

yepp..thats the question! i wasnt too sure on how to use the circumflex looking like thing!

OpenStudy (mertsj):

Get it equal to 0 and factor it.

OpenStudy (mertsj):

\[3(1+\tan ^2x)-6=8\tan x\]

OpenStudy (mertsj):

\[3\tan ^2x-8\tan x-3=0\]

OpenStudy (mertsj):

\[(3\tan x+1)(\tan x-3)=0\]

OpenStudy (mertsj):

Take it from there.

OpenStudy (anonymous):

thanks!

OpenStudy (mertsj):

yw